Which of the ordered pairs
(–7, 2), (1, 0), (–3, –1), (–23, 6)
are solutions for the equation x + 4y = 1?
Thanks so much for your valuable time!! No book.

Answer by ankor@dixie-net.com(22740)

Which of the ordered pairs
(–7, 2), (1, 0), (–3, –1), (–23, 6)
:
are solutions for the equation x + 4y = 1?
:
Each ordered pair represents the values of x & y.
:
Substitute the values for x & y and see if the equation is true;
:
The first pair: (-7,2), substitute these value for x & y in x + 4y = 1
:
-7 + 4(2) =
-7 + 8 = 1, so the 1st pair is a solution
:
The 2nd pair: (1,0)
:
1 + 4(0) = 1
1 + 0 = 1 so the 2nd pair is also a solution
:
The 3d pair: (-3,-1)
:
-3 + 4(-1) = 1
-3 - 4 = -7 so this is not a solution
:
The 4th pair: (-23,6)
: -23 + 4(6) = 1
-23 + 24 = 1 so this is a solution also